OpenStudy (anonymous):

2(6-3)+8/2 show work

OpenStudy (anonymous):

2(6-3)+8/2 = 12-6+8/2 = 12-6+4 = 10 --> is this the ans u r looking 4?

OpenStudy (anonymous):

the answer would be 7... you do the order of operations. Please Excuse My Dear Ant Sally

OpenStudy (anonymous):

i don't get it... why?

OpenStudy (anonymous):

You do paranthesis first... so it would be 2 times 3, not multiplication where you distrbute it where it would be 12-6

OpenStudy (anonymous):

oh... thanks!